Required information [The following information applies to the questions displayed below.] Sweeten Company had no jobs in pro1. What was the companys plantwide predetermined overhead rate? (Round your answer to 2 decimal places.) Predetermined overh2. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calcula3. What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.) Total manufacturing cos4. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ5. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.) Total manufacturing cos6. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establi8. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold9. What were the companys predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your10. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do n11. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.) Unit product cost13. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final ans14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establ15. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold
Answer #1

1) Overhead rate = (2500*2.8+1500*3.6+30600)/4000 = 10.75 per hour

2) Manufacturing overhead applied

Job P Job Q
Manufacturing overhead applied 5100*10.75 = 54825 48375

3) Total manufacturing cost of Job P

Job P
Direct material 27000
Direct labor 32200
Overhead 54825
Total 114025

4) Unit product cost = 114025/20 = 5701

Note: Please post whole question in the group of 4-4 sub parts as per HOMEWORKLIB RULES